OpenStudy (anonymous):

Graph the inequalities 4x – 5y –10 x + 3y –12

OpenStudy (anonymous):

do you have a greater than sign or a less than sign?

OpenStudy (anonymous):

4x – 5y <_ –10 x + 3y _> –12

OpenStudy (mertsj):

Can you solve them for y?

OpenStudy (anonymous):

Thats what I need help with

OpenStudy (mertsj):

\[4x-5y \le-10\]

OpenStudy (mertsj):

Subtract 4x from both sides

OpenStudy (mertsj):

\[-5y \le-4x-10\]

OpenStudy (anonymous):

Yea I did that

OpenStudy (mertsj):

Now divide both sides by -5 and don't forget to reverse the inequality symbol.

OpenStudy (anonymous):

Then divide -5 does the sign flip?

OpenStudy (mertsj):

\[y \ge\frac{4}{5}x+2\]

OpenStudy (mertsj):

Now graph it. Mark 2 on the y axis. From that point, go up 4 and right 5

OpenStudy (mertsj):



OpenStudy (mertsj):

Ok. Are you ready for the second one?

OpenStudy (mertsj):

\[x+3y \ge-12\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Subtract x from both sides.

OpenStudy (mertsj):

\[3y \ge-x-12\]

OpenStudy (mertsj):

Now divide both sides by 3

OpenStudy (mertsj):

\[y \ge-\frac{1}{3}x-4\]

OpenStudy (mertsj):

Now mark -4 on the y axis and then go down 1 and right 3.

OpenStudy (mertsj):